OpenStudy (anonymous):

If you are finding the line perpendicular to x + 3y = 5, what is the slope of the perpendicular line? A. -3/5 B. 1/3 C. 3 D. -3

OpenStudy (anonymous):

Do you really need help on this?

OpenStudy (anonymous):

-.-

OpenStudy (anonymous):

Can you change that line into slope-intercept form for me please

OpenStudy (anonymous):

Abii are you there mam

OpenStudy (anonymous):

now i am

OpenStudy (anonymous):

Okay. Do it

OpenStudy (anonymous):

is it b ?

OpenStudy (anonymous):

Abii que haces mija

OpenStudy (anonymous):

Yes, but are you just guessing.

OpenStudy (anonymous):

It seems like you get it, but you're still asking the questions.

OpenStudy (anonymous):

no im not o:

OpenStudy (anonymous):

Wait

OpenStudy (anonymous):

Nvm

OpenStudy (anonymous):

Okay. So change it into slope-intercept form so we can do this

OpenStudy (anonymous):

y=mx+b c'mon you know this

OpenStudy (anonymous):

this is too hard D;

OpenStudy (anonymous):

Y=MX+B x+3y=5

OpenStudy (anonymous):

What do you have to do to make it look like y=mx+b?

OpenStudy (anonymous):

What can we manipulate to get the y on one side and the slope and the y-intercept on the other.

OpenStudy (anonymous):

3 ?

OpenStudy (anonymous):

You subtract x so you get 3y=-x+5 What do we do now to make y has a coefficient of 1?

OpenStudy (anonymous):

have*

OpenStudy (anonymous):

umm.... you subtract 3 from both sides ?

OpenStudy (anonymous):

No. You can't subtract because it's being multiplied to the y. So the opposite of multiplication is division. We're gonna divide by 3

OpenStudy (anonymous):

So you end up with y= -1/3x+5/3

OpenStudy (anonymous):

Okay so now we can find what the slope has to be for a perpendicular line All we're worrying about is the slope which is -1/3 In order for the other line to be perpendicular, the slope has to be the opposite reciprocal Meaning you flip the fraction, and you change the sign. So if it's negative, it turns positive, and vice versa. For Example: If the slope was 2 For it to be perpendicular you would need a slope of -1/2 Okay now do it with the slope that was given to us. What slope would we need to be perpendicular to -1/3

OpenStudy (anonymous):

... Did you do it?

OpenStudy (anonymous):

i have noo idea what im doing /.<

OpenStudy (anonymous):

WE have the slope. I just told you it's -1/3. Take the opposite reciprocal of that and that's the answer. I gave you an example and instructions on how to do it already.

OpenStudy (anonymous):

okay smart retrice-.- thank you for your help

OpenStudy (anonymous):

Alright. You got it?

OpenStudy (anonymous):

*retrice

OpenStudy (anonymous):

Yes. I got it. Retrice

OpenStudy (anonymous):

yes i got it

OpenStudy (anonymous):

What did you get?

OpenStudy (anonymous):

the whole conception

OpenStudy (anonymous):

What's your answer.

OpenStudy (anonymous):

b

OpenStudy (anonymous):

*concept

OpenStudy (anonymous):

Conception is getting pregnant.

OpenStudy (anonymous):

No It's not B

OpenStudy (anonymous):

ohh lmfao and yes it iss ??

OpenStudy (anonymous):

Opposite Reciprocal. Let's do this TWO-STEP process one step at a time, because it's a bit too complicated for you. TAke the slope: -1/3 First flip the fraction

OpenStudy (anonymous):

its okay i got it teacher -.-

OpenStudy (anonymous):

What is it

OpenStudy (anonymous):

your making my life complicated right now -.-

OpenStudy (anonymous):

Flip the fraction. Flip IT. Just like you flip 1/2 into 2/1=2 Or 1/3 into 3/1 or 5/4 into 4/5

OpenStudy (anonymous):

3/1

OpenStudy (anonymous):

Wow you did the two steps already. That's what reciprocal means. Flip the fraction. So the reciiprocal of -1/3 is -3/1=-3 and the other step is to change the signs (opposite). So the answer is 3/1=3 Finally

OpenStudy (anonymous):

am i done now teacher ? can i go to recess ?

OpenStudy (anonymous):

I gave you the answer didn't I? SO NOW WHENEVER SOMEONE TELLS YOU TO TAKE THE OPPOSITE RECIPROCAL YOU DO IT.

OpenStudy (anonymous):

Sleep tight